# Towards the Usage of Window Counting Constraints

### v1

@@ Introduction
“The system plays action $act$ at least $k$ times out of $l$ of its own moves.”
@@
This work is an extension of [FF24]. The extended version complements the content of the original one in two main areas: For once, the synthesis approach in [FF24] only considered safety games, while this work extends the game types to reachability, Büchi, co-Büchi and parity games. Secondly, this work adds the possibility of iterating over constraints of both the form “The system plays action $act$ at least $k$ times out of $l$ of its own moves.” and “The system plays action $act$ at most $k$ times out of $l$ of its own moves.” in the same synthesis procedure, while the previous version of the approach considered one of the constraint types to be fixed. Additionally, this article is enriched with more extensive explanations and examples.
@@ 1 Related Work
2 Games with Counting Constraints
@@
A game $G=(A,\text{Win})$ consists of a game graph $A$ and a winning condition Win like a reachability, safety, Büch, co-Büchi or parity condition. The winning condition consist of a set of (infinite) sequences of states of $G$. $EGO$ wins if it has a strategy that guarantees to only plays in Win. The environment, on the other hand, wins if a play not in Win is played. Hence, each play of a two-player safety game always has exactly one winner and one loser. Games with this property are called zero sum games.
@@
  1. 1.
    For $p=EGO$, $CC_{max}(EGO,a,k,l)$ is defined as the abbreviation for “The player $EGO$ plays $a$ at most $k$ times out of $l$ of its own turns.”. $CC_{max}(EGO,a,k,l)$ is satisfied on $\pi$, if for all $i\in\mathbb{N}_{0}$ holds |$|\{\sigma_{2m}\,|\,\sigma_{2m}\models a,\,i\leq m\leq i+l\}|\leq k$.
  2. 2.
    For $p=EGO$, $CC_{min}(EGO,a,k,l)$ is defined as the abbreviation for “The player $EGO$ plays $a$ at least $k$ times out of $l$ of its own turns.”. $CC_{min}(EGO,a,k,l)$ is satisfied on $\pi$, if for all $i\in\mathbb{N}_{0}$ holds |$|\{\sigma_{2m}\,|\,\sigma_{2m}\models a,\,i\leq m\leq i+l\}|\geq k$.
  3. 3.
    For $p=ALTER$, $CC_{max}(ALTER,a,k,l)$ is defined as the abbreviation for “The player $ALTER$ plays $a$ at most $k$ times out of $l$ of its own turns.”. $CC_{max}(ALTER,a,k,l)$ is satisfied on $\pi$, if for all $i\in\mathbb{N}_{0}$ holds |$|\{\sigma_{2m+1}\,|\,\sigma_{2m+1}\models a,\,i\leq m\leq i+l\}|\leq k$.
  4. 4.
    For $p=ALTER$, $CC_{min}(ALTER,a,k,l)$ is defined as the abbreviation for “The player $ALTER$ plays $a$ at least $k$ times out of $l$ of its own turns.”. $CC_{min}(ALTER,a,k,l)$ is satisfied on $\pi$, if for all $i\in\mathbb{N}_{0}$ holds |$|\{\sigma_{2m+1}\,|\,\sigma_{2m+1}\models a,\,i\leq m\leq i+l\}|\geq k$.
@@
The prefix $pre=(1,\{x\},4,\{a\},1,\{x\},4,\{x\},1,\{x\},4)$ of another play on the shown graph is not satisfying “The player $EGO$ plays $y$ at least $1$ times out of $5$ of its own turns.”, since no matter what $EGO$ does in its next move (even if it would be allowed to change the game graph for the next turns), it is not possible to extend $pre$ to a play that satisfies the constraint.
11112222333344445555yaxx, ybcyax
Figure 1: Small game graph for a two-player game. Circles represent states controlled by $EGO$, diamond-shaped states are controlled by $ALTER$. Transitions are equipped with the actions a player plays when deciding for a transition. For example, $EGO$ can choose in state 3 to play $x$ and $y$ in one turn (leading to state 4) or just play $x$ (and not $y$), which leads to state 2.
$\pi=$$1,$ $\{\mathbf{y}\},$ $2,$ $\{a\},$ $3,$ $\{\mathbf{x}\},$ $2,$ $\{a\},$ $3,$ $\{\mathbf{x},\mathbf{y}\},$ $4,$ $\{a\},$ $1,$ $\{\mathbf{y}\},$ $2,$ $\{a\},$ $3,$ $\{\mathbf{x}\},$ $2,$ $\{a\},$ $3,$ $\{\mathbf{x},\mathbf{y}\},$dots $x$ in at least 2 out of 4 turns of $EGO$$x$ in at least 2 out of 4 turns of $EGO$$x$ in at least 2 out of 4 turns of $EGO$$y$ in at most 2 out of 3 turns of $EGO$$y$ in at most 2 out of 3 turns of $EGO$$y$ in at most 2 out of 3 turns of $EGO$$y$ in at most 2 out of 3 turns of $EGO$
Figure 2: Illustration of the “sliding window”-property of window counting constraints: Each constraint needs to be fulfilled at each part of a play. Actions in $\pi$ played by $EGO$ are highlighted in bold. $\pi$ is a prefix of a play in the game graph shown in Figure 1.
In a (zero-sum) game with counting constraints, the $EGO$ player needs to satisfy all of its counting constraints in order to win the game. We assume that the player $ALTER$ also plays according to its counting constraints. This is an assumption on the rationality of $ALTER$: $ALTER$ still plays antagonistic against $EGO$, but not if that means to violate the own specifications. For now, we only consider games in which $ALTER$ can always find a way to satisfy its counting constraints. In particular, $EGO$ cannot force $ALTER$ into a constraint violation.
@@
  • $CC$ is a finite sets of counting constraints.
  • $ALTER$ cannot be forced into constraint violations: For each prefix $\pi(n)=\pi_{0}\sigma_{0}\pi_{1}\dots\pi_{n-1}$$\sigma_{n-1}\pi_{n}$, $n\in 2\mathbb{N}+1$, of a play on $G$ that satisfies all counting constraints of $ALTER$, there exists $(\pi_{n},a,\pi_{n+1})\in\rightarrow$, such that $\pi_{0}\sigma_{0}\pi_{1}\dots\pi_{n-1}\sigma_{n-1}\pi_{n}a\pi_{n+1}$ is also a prefix of a play on $G$ that satisfies the counting constraints of $ALTER$.
@@
The fourth property of a game with counting constraints seems counter-intuitive, since it restricts the considered games to those in which $ALTER$ can basically ensure in each step to satisfy all of its constraints without any need of planning ahead. The main reason for introducing this property is as follows: We are not interested in plays that are only won by $EGO$ due to forcing $ALTER$ to loose due to violated counting constraints. Dismissing the property would lead to situations like depicted in Figure 3: The $EGO$-player does not have any counting constraints to obey but needs to avoid an unsafe state (colored in gray). $ALTER$ has the counting constraint $ALTER$ plays $b$ at least 1 time in 1 turn.” (hence, $ALTER$ needs to play $b$ in each of its turns). In the initial state, choosing to enter the upper sub-graph lets $EGO$ win. However, taking the transitions to the lower subgraph and consequently visiting an unsafe state would also lead to a win for $EGO$, since $ALTER$ is violating its constraint in the resulting play. This problem of winning by falsifying the assumptions, but not leading to intended behavior of the player is well-known [MPS19], [BEJK14].
@@
Figure 3: Graph for a safety game that cannot be used in a game with the counting constraint $ALTER$ plays $b$ at least 1 time in 1 turn.”. Circles represent states controlled by $EGO$, diamond-shaped states are controlled by $ALTER$. The state colored in gray marks an unsafe state. When the lower subgraph is entered, $A$ cannot fulfill its counting constraint, violating the fourth property in Definition 2.
@@
3 Incremental Synthesis with Counting Constraints
The key advantage of counting constraints for synthesis is their monotonicity property: If $EGO$ has a strategy, such that $EGO$ plays an action $a$ at most $k$ times in $l$ turns (i.e. the strategy satisfies $CC_{max}(EGO,a,k,l)$), then $EGO$ also plays $a$ at most $k$ times in $l-1$ turns (i.e. the strategy satisfies $CC_{max}(EGO,a,k,l-1)$). In other words: The existence of a winning strategy for a game with counting constraint $CC_{max}(EGO,a,k,l-1)$ is a necessary condition for the winning strategy for a game with $CC_{max}(EGO,a,k,l)$. Moreover, only a strategy that fulfills $CC_{max}(EGO,a,k,l-1)$ can also fulfill $CC_{max}(EGO,a,k,l)$. From an algorithmic perspective, it is more favorable to search for strategies that satisfy $CC_{max}(EGO,a,k,l-1)$ then for strategies that satisfy $CC_{max}(EGO,a,k,l)$, since the graph that encodes the first (shorter) constraint is smaller than the one that encodes the latter (longer) constraint. Intuitively, this is caused by more memory that is needed for remembering the last $l$ own turns instead of only $l-1$ turns. The synthesis idea is related to the incremental approach used by synthesis with antichains [FJR09].
For a counting constraint of the form $CC_{min}(EGO,a,k,l-1)$ ($EGO$ plays action $a$ at least $k$ times out of $l-1$ of its turns”), we can conduct that if a strategy fulfills the constraint, it automatically also fulfills the larger constraint $CC_{min}(EGO,a,k,l)$. Hence, if we already have a strategy that fulfills $CC_{min}(EGO,a,k,l-1)$, it is needless to do the more challenging search for a strategy that fulfills $CC_{min}(EGO,a,k,l)$.
@@ 3 Incremental Synthesis with Counting Constraints
The player $EGO$ plays $a$
at most $k$ times
out of $l$ of its own turns.
The player $EGO$ plays $a$
at most $l-k+1$ times
out of $l+1$ of its own turns.
The player $EGO$ plays $\neg a$
at least $k$ times
out of $l$ of its own turns.
The player $EGO$ plays $\neg a$
at least $k$ times
out of $l+1$ of its own turns.
Combining Theorem 1 and Lemma 2 yields for later increments: Instead of directly considering a counting constraint of the form “The player $EGO$ plays $a$ at most $k$ times out of $l$ of its own turns.”, we can start with “The player $EGO$ plays $\neg a$ at least $l-k$ times out of $l-k$ of its own turns.”, incrementally increase the constraint length to “The player $EGO$ plays $\neg a$ at least $l-k$ times out of $l$ of its own turns.”, which translated back to the original counting constraint.
@@
Figure 4: Snippet of a situation graph for the game from Example 2. A state (“situation”) of the game graph is shown as a combination of a state from the original game graph (Figure 1) with a history that allows to check counting constraint fulfillment in the situation. For better readability, the history is written as table with the name of the counting constraints as header: $C_{1}=$ “The player $EGO$ plays $x$ at least $2$ times out of $4$ of its own turns.”, $C_{2}=$ “The player $EGO$ plays $y$ at least $1$ times out of $5$ of its own turns.”, $C_{3}=$ “The player $EGO$ plays $y$ at most $2$ times out of $3$ of its own turns.”, $C_{4}=$ “The player $ALTER$ plays $a\vee b$ at least $1$ time out of $2$ of its own turns.”.
@@
For each counting constraint $CC_{m}(p,a,k,l)$ with $m\in\{min,max\}$, $p\in\{EGO,ALTER\}$, $a$ being a logical formula over $(\Sigma_{p})$ as defined in Definition 2 , $k,l\in\mathbb{N}$, $k\leq l$ define a mapping
@@
$w_{i}=\biggl{\{}\begin{aligned} h_{CC_{i}}(v_{i},act),&\text{ if }s\text{ is controlled by the player that has to satisfy }CC_{i}\\ v_{i},&\text{ else.}\end{aligned}$
@@
  • initial state being the situation $s_{init}=(s_{0},H_{init})$ with all entries in $H_{init}$ being $none$,
@@
  • $\text{Win}^{\prime}$ the winning condition for $EGO$: Extend the winning condition $\text{Win}\subseteq S^{\omega}$ in $G$ to the situation graph by defining $\text{Win}^{\prime}=\{(s_{i},H_{i})_{i\in\mathbb{N}}\,|\,(s_{i})_{i\in\mathbb{N}}\in\text{Win}\text{ and for each }i\in\mathbb{N}\text{ holds: }(s_{i},H_{i})\text{ satisfies all counting constraints from $CC$ on }EGO\}$
@@
The general idea is to start with a rather small game by using counting constraints of small lengths and increment over the length. In each increment, a part of the corresponding situation graph is constructed and analyzed and knowledge that can be reused in following increments is identified. This knowledge is determining which parts of the situation graph for the next increment needs to be constructed and which parts can be omitted, relying on 1. To this end, we introduce extensions of situations in order co compare situations from one increment with situations in another increment. {defi}[Extensions of situations] Let $(s,H)$ be a situation with counting constraints $CC$, $H=(H_{1},\dots,H_{n})$, $H_{i}$ representing the history for counting constraint $CC_{i}\in CC$. Let $I$ be the set of indices of counting constraints of the form $CC_{min}(EGO,a,k,l)$. Similarly, let $(\tilde{s},\tilde{H})$ be another situation with counting constraints $\widetilde{CC}$, $\tilde{H}=(\tilde{H}_{1},\dots,\tilde{H}_{m})$, $\tilde{H}_{i}$ representing the history for counting constraint $\widetilde{CC}_{i}\in\widetilde{CC}$. $(\tilde{s},\tilde{H})$ is an extension of $(s,H)$, if $s=\tilde{s}$, $n=m$ and if it exists a permutation $\pi$ on $\{1,\dots,n\}$, such that the following properties all hold:
@@
  • $(s,H)$ and $(\tilde{s},\tilde{H})$ do not differ in all other counting constraints and history vectors (except for the ordering): $\forall i\notin I:\,\widetilde{CC}_{\pi(i)}=CC_{i}\text{ and }\tilde{H}_{\pi(i)}=H_{i}$
@@
Figure 5: Examples for situations being extensions of other situations. The notation is the same as in Example 3 and Figure 4.
@@
Figure 6: Overview of the incremental synthesis algorithm for a game with counting constraints $G=(A,\text{Win},CC)$ with increments over constraints of the form $CC_{min}(EGO,a,k,l)$
@@
  • Situations that violate any constraint of $ALTER$ are not added to the situation graph. For any situation that satisfies all constraints of player $A$ and is controlled by $ALTER$, $ALTER$ can choose an action as next turn, such that the reached situation also satisfies all of its counting constraints (as a consequence of the last property in Definition 2 and Theorem 1). Hence, all relevant behavior of $ALTER$ is still captured when omitting situations that violate $ALTER$-constraints.
@@
Figure 7: Illustration on how the situation graph is modified in the incremental synthesis over counting constraint length as depicted in Figure 6. Circles represent situations controlled by $EGO$, diamond-shaped situations are controlled by $ALTER$. Situations colored in gray represent extensions of situations in the winning region of a previous increment. Situations marked with doubled lines are considered to have a history that violates a constraint of $EGO$. Similarly, histories of situations with dashed lines are violating a constraint of $ALTER$.
@@
We define two sinks $W$ and $L$ that are incorporated into the situation graph for each winning condition and then go into specific adaptions for a selected set of winning conditions (safety, reachability, Büchi, co-Büchi, parity). For staying in a two-player game, both sinks consists of one state controlled by $EGO$ and one state controlled by $ALTER$ with transitions between the two states in both directions. The history-part of the states and the actions played when choosing one of the transitions are irrelevant in the following process. Denote the two states of $W$ with $(w_{E},\boldsymbol{\cdot})$ (controlled by $EGO$) and $(w_{A},\boldsymbol{\cdot})$ (controlled by $ALTER$) with $\boldsymbol{\cdot}$ being a placeholder for any history and $w_{E}$, $w_{A}$ symbolizing states that are not already in the original game graph $A$. Likewise, denote the states in $L$ as $(l_{E},\boldsymbol{\cdot})$ and $(l_{A},\boldsymbol{\cdot})$. The sinks are integrated into the (modified) situation graph from the previous step as follows:
@@
Denote the resulting graph with sinks as $Sit_{\mathcal{C}}=(\tilde{S},s_{init},\tilde{S}_{EGO},\tilde{S}_{ALTER},\tilde{\hookrightarrow})$ with $\tilde{S}$ being all situations in the modified situation graph of the previous step plus the four sink states, $s_{init}$ the initial state of the situation graph as in Definition 4, $\tilde{S}_{EGO}$, $\tilde{S}_{ALTER}$ forming a disjoint union of $\tilde{S}$ as situations controlled by $EGO$ and $ALTER$, respectively, and $\tilde{\hookrightarrow}$ denoting the transitions in the graph with sinks. With a slightly abuse of notation, we call the resulting graph $Sit_{\mathcal{C}}$ with sinks again situation graph. Figure 7(c) illustrates the integration of the two sinks $W$ and $L$.
Safety winning condition $Win=\mathtt{Safe}(R)$: With $\tilde{R}:=\{(s,H)\in\tilde{S}\,|\,s\in R\cup\{(w_{E},w_{A})\}\}$ and winning condition $\widetilde{\text{Win}}=\mathtt{Safe}(\tilde{R})$ is the game $(Sit_{\mathcal{C}},\widetilde{\text{Win}})$ now solvable with any synthesis algorithm that is also applicable for solving the original game $(A,\text{Win})$.
Büchi winning condition $\text{Win}=\mathtt{B\ddot{u}chi}(R)$: Similar to the safety winning condition case, define $\tilde{R}:=\{(s,H)\in\tilde{S}\,|\,s\in R\cup\{(w_{E},w_{A})\}\}$. With $\widetilde{\text{Win}}:=\mathtt{B\ddot{u}chi}(\tilde{R})$, the game $(Sit_{\mathcal{C}},\widetilde{\text{Win}})$ is a Büchi game that can be solved by any synthesis algorithm that can solve $(A,\text{Win})$. All $\pi\in\mathtt{B\ddot{u}chi}(\tilde{R})$ do not violate any counting constraint in $\mathcal{C}$, since each constraint violation automatically leads to sink $L$, hence any play violating any of the constraint in any visited situation automatically visits $(l_{E},\boldsymbol{\cdot})$ and $(l_{A},\boldsymbol{\cdot})\notin\tilde{R}$ infinitely often.
Co-Büchi winning condition $\text{Win}=\mathtt{coB\ddot{u}chi}(R)$: With $\tilde{R}:=\{(s,H)\in\tilde{S}\,|\,s\in R\cup\{(w_{E},w_{A})\}\}$ and $\widetilde{\text{Win}}:=\mathtt{coB\ddot{u}chi}(\tilde{R})$, the same argumentation as for Büchi winning conditions holds.
@@
as coloring on $Sit_{\mathcal{C}}$. Situations (except for the new sink-locations) inherit their color from their state in $S$. Whenever a play enters the winning sink $W$, $EGO$ will win the play, since each situation in $W$ is visited infinitely often and is colored with 0, being the minimal coloring and even. Similarly, if a play enters the sink $L$, $EGO$ is going to loose, since $L$ will never be left, hence only finitely many visited situations are not in $L$ and all situations in $L$ are colored with the uneven color 1.
@@
  2. 2.
    Define $\tilde{R}:=\{(s,H)\in\mathcal{W}\,|\,s\in R\cup\{w_{E},w_{A}\}\}$. With $\widetilde{\text{Win}}:=\mathtt{Reach}(\tilde{R})$, the game $(\mathcal{W},\widetilde{\text{Win}})$ becomes a two-player reachability game with winning plays containing at least one situation with state in $R$ and not violating any counting constraint of $EGO$.
With those modifications, the (adapted) situation graph with (adapted) winning condition becomes a game $G_{\mathcal{C}}(Sit_{\mathcal{C}},\widetilde{\text{Win}})$ of the same type as the original game $G=(A,\text{Win})$, such that a synthesis algorithm of the readers choice can be applied to determine the winning region. Note that the construction of $\widetilde{\text{Win}}$ is in this form only possible since we only increment over counting constraints of the form $CC_{min}(EGO,a,k,l)$ (with constraints of the form $CC_{max}(EGO,a,k,l)$ optionally translated by using Lemma 2). A key property that we are using here is that extensions of situations in the winning region of one increment are again in the winning region. If incrementing would go directly over $CC_{max}(EGO,a,k,l)$ with increasing constraint length, this property would not hold.
Determine the winning region of $G_{\mathcal{C}}(Sit_{\mathcal{C}},\widetilde{\text{Win}})$
In this step, the winning region of $G_{\mathcal{C}}=(Sit_{\mathcal{C}},\widetilde{Win})$ with the constructions of the previous step is calculated by any synthesis algorithm that can calculate the winning region of the original game $G=(A,\text{Win})$ without window counting constraints. The calculated winning region serves as input for following increments, especially for constructing the next situation graph. In fact, this information on the winning region is the key enabler to speed up the next increment. Note that the calculated winning region in this step is in general only a subset of the winning region of $(A,\text{Win},\mathcal{C})$, as a price for not extending the situation graph in situations that are extensions of the winning region of previous increments. However, this is not a problem as long as the overall goal is to find a winning strategy for $EGO$ and not the full winning region. We can just observe that $EGO$ will satisfy shorter (and hence more restrictive) counting constraints than necessary when using the presented incremental approach.

### v2

@@ Introduction
“The system plays action actact at least $k$ times out of $l$ of its own moves.”
@@
This work is an extension of [FF24]. The extended version complements the content of the original one in two main areas: For once, the synthesis approach in [FF24] only considered safety games, while this work extends the game types to reachability, Büchi, co-Büchi and parity games. Secondly, this work adds the possibility of iterating over constraints of both the form “The system plays action actact at least $k$ times out of $l$ of its own moves.” and “The system plays action actact at most $k$ times out of $l$ of its own moves.” in the same synthesis procedure, while the previous version of the approach considered one of the constraint types to be fixed. Additionally, this article is enriched with more extensive explanations and examples.
@@ 1. Related Work
2. Games with Counting Constraints
@@
A game $G=(A,\text{Win})$ consists of a game graph $A$ and a winning condition Win like a reachability, safety, Büch, co-Büchi or parity condition. The winning condition consist of a set of (infinite) sequences of states of $G$. EGOEGO wins if it has a strategy that guarantees to only plays in Win. The environment, on the other hand, wins if a play not in Win is played. Hence, each play of a two-player safety game always has exactly one winner and one loser. Games with this property are called zero sum games.
@@
  1. (1)
    For $p=EGO$, $CC_{max}(EGO,a,k,l)$ is defined as the abbreviation for “The player $EGO$ plays $a$ at most $k$ times out of $l$ of its own turns.”. $CC_{max}(EGO,a,k,l)$ is satisfied on $\pi$, if for all $i\in\mathbb{N}_{0}$ holds $|\{\sigma_{2m}\,|\,\sigma_{2m}\models a,\,i\leq m\leq i+l\}|\leq k$.
  2. (2)
    For $p=EGO$, $CC_{min}(EGO,a,k,l)$ is defined as the abbreviation for “The player $EGO$ plays $a$ at least $k$ times out of $l$ of its own turns.”. $CC_{min}(EGO,a,k,l)$ is satisfied on $\pi$, if for all $i\in\mathbb{N}_{0}$ holds $|\{\sigma_{2m}\,|\,\sigma_{2m}\models a,\,i\leq m\leq i+l\}|\geq k$.
  3. (3)
    For $p=ALTER$, $CC_{max}(ALTER,a,k,l)$ is defined as the abbreviation for “The player $ALTER$ plays $a$ at most $k$ times out of $l$ of its own turns.”. $CC_{max}(ALTER,a,k,l)$ is satisfied on $\pi$, if for all $i\in\mathbb{N}_{0}$ holds $|\{\sigma_{2m+1}\,|\,\sigma_{2m+1}\models a,\,i\leq m\leq i+l\}|\leq k$.
  4. (4)
    For $p=ALTER$, $CC_{min}(ALTER,a,k,l)$ is defined as the abbreviation for “The player $ALTER$ plays $a$ at least $k$ times out of $l$ of its own turns.”. $CC_{min}(ALTER,a,k,l)$ is satisfied on $\pi$, if for all $i\in\mathbb{N}_{0}$ holds $|\{\sigma_{2m+1}\,|\,\sigma_{2m+1}\models a,\,i\leq m\leq i+l\}|\geq k$.
@@
The prefix $pre=(1,\{x\},4,\{a\},1,\{x\},4,\{x\},1,\{x\},4)$ of another play on the shown graph is not satisfying “The player $EGO$ plays $y$ at least $1$ times out of $5$ of its own turns.”, since no matter what $EGO$ does in its next move (even if it would be allowed to change the game graph for the next turns), it is not possible to extend prepre to a play that satisfies the constraint.
1122334455yaxx, ybcyax
Figure 1. Small game graph for a two-player game. Circles represent states controlled by $EGO$, diamond-shaped states are controlled by $ALTER$. Transitions are equipped with the actions a player plays when deciding for a transition. For example, EGOEGO can choose in state 3 to play $x$ and $y$ in one turn (leading to state 4) or just play $x$ (and not $y$), which leads to state 2.
$\pi=$$1,$ $\{\mathbf{y}\},$ $2,$ $\{a\},$ $3,$ $\{\mathbf{x}\},$ $2,$ $\{a\},$ $3,$ $\{\mathbf{x},\mathbf{y}\},$ $4,$ $\{a\},$ $1,$ $\{\mathbf{y}\},$ $2,$ $\{a\},$ $3,$ $\{\mathbf{x}\},$ $2,$ $\{a\},$ $3,$ $\{\mathbf{x},\mathbf{y}\},$ $\dots$ $x$ in at least 2 out of 4 turns of EGOEGOxx in at least 2 out of 4 turns of EGOEGOxx in at least 2 out of 4 turns of EGOEGOyy in at most 2 out of 3 turns of EGOEGOyy in at most 2 out of 3 turns of EGOEGOyy in at most 2 out of 3 turns of EGOEGOyy in at most 2 out of 3 turns of EGOEGO
Figure 2. Illustration of the “sliding window”-property of window counting constraints: Each constraint needs to be fulfilled at each part of a play. Actions in $\pi$ played by EGOEGO are highlighted in bold. $\pi$ is a prefix of a play in the game graph shown in Figure 1.
In a (zero-sum) game with counting constraints, the $EGO$ player needs to satisfy all of its counting constraints in order to win the game. We assume that the player $ALTER$ also plays according to its counting constraints. This is an assumption on the rationality of $ALTER$: ALTERALTER still plays antagonistic against $EGO$, but not if that means to violate the own specifications. For now, we only consider games in which $ALTER$ can always find a way to satisfy its counting constraints. In particular, $EGO$ cannot force $ALTER$ into a constraint violation.
@@
  • CCCC is a finite sets of counting constraints.
  • ALTERALTER cannot be forced into constraint violations: For each prefix $\pi(n)=\pi_{0}\sigma_{0}\pi_{1}\dots\pi_{n-1}$$\sigma_{n-1}\pi_{n}$, $n\in 2\mathbb{N}+1$, of a play on $G$ that satisfies all counting constraints of $ALTER$, there exists $(\pi_{n},a,\pi_{n+1})\in\rightarrow$, such that $\pi_{0}\sigma_{0}\pi_{1}\dots\pi_{n-1}\sigma_{n-1}\pi_{n}a\pi_{n+1}$ is also a prefix of a play on $G$ that satisfies the counting constraints of $ALTER$.
@@
The fourth property of a game with counting constraints seems counter-intuitive, since it restricts the considered games to those in which $ALTER$ can basically ensure in each step to satisfy all of its constraints without any need of planning ahead. The main reason for introducing this property is as follows: We are not interested in plays that are only won by $EGO$ due to forcing $ALTER$ to loose due to violated counting constraints. Dismissing the property would lead to situations like depicted in Figure 3: The $EGO$-player does not have any counting constraints to obey but needs to avoid an unsafe state (colored in gray). $ALTER$ has the counting constraint ALTERALTER plays $b$ at least 1 time in 1 turn.” (hence, $ALTER$ needs to play $b$ in each of its turns). In the initial state, choosing to enter the upper sub-graph lets $EGO$ win. However, taking the transitions to the lower subgraph and consequently visiting an unsafe state would also lead to a win for $EGO$, since $ALTER$ is violating its constraint in the resulting play. This problem of winning by falsifying the assumptions, but not leading to intended behavior of the player is well-known [MPS19], [BEJK14].
@@
Figure 3. Graph for a safety game that cannot be used in a game with the counting constraint ALTERALTER plays $b$ at least 1 time in 1 turn.”. Circles represent states controlled by $EGO$, diamond-shaped states are controlled by $ALTER$. The state colored in gray marks an unsafe state. When the lower subgraph is entered, AA cannot fulfill its counting constraint, violating the fourth property in Definition 2.
@@
3. Incremental Synthesis with Counting Constraints
The key advantage of counting constraints for synthesis is their monotonicity property: If $EGO$ has a strategy, such that $EGO$ plays an action $a$ at most $k$ times in $l$ turns (i.e. the strategy satisfies $CC_{max}(EGO,a,k,l)$), then $EGO$ also plays $a$ at most $k$ times in l1l-1 turns (i.e. the strategy satisfies $CC_{max}(EGO,a,k,l-1)$). In other words: The existence of a winning strategy for a game with counting constraint $CC_{max}(EGO,a,k,l-1)$ is a necessary condition for the winning strategy for a game with $CC_{max}(EGO,a,k,l)$. Moreover, only a strategy that fulfills $CC_{max}(EGO,a,k,l-1)$ can also fulfill $CC_{max}(EGO,a,k,l)$. From an algorithmic perspective, it is more favorable to search for strategies that satisfy $CC_{max}(EGO,a,k,l-1)$ then for strategies that satisfy $CC_{max}(EGO,a,k,l)$, since the graph that encodes the first (shorter) constraint is smaller than the one that encodes the latter (longer) constraint. Intuitively, this is caused by more memory that is needed for remembering the last $l$ own turns instead of only l1l-1 turns. The synthesis idea is related to the incremental approach used by synthesis with antichains [FJR09].
For a counting constraint of the form $CC_{min}(EGO,a,k,l-1)$ (EGOEGO plays action $a$ at least $k$ times out of l1l-1 of its turns”), we can conduct that if a strategy fulfills the constraint, it automatically also fulfills the larger constraint $CC_{min}(EGO,a,k,l)$. Hence, if we already have a strategy that fulfills $CC_{min}(EGO,a,k,l-1)$, it is needless to do the more challenging search for a strategy that fulfills $CC_{min}(EGO,a,k,l)$.
@@ 3. Incremental Synthesis with Counting Constraints
The player $EGO$ plays $a$ at most $k$ times out of $l$ of its own turns. The player $EGO$ plays $a$ at most lk+1l-k+1 times out of $l+1$ of its own turns. The player $EGO$ plays $\neg a$ at least $k$ times out of $l$ of its own turns. The player $EGO$ plays $\neg a$ at least $k$ times out of $l+1$ of its own turns.
Combining Theorem 1 and Lemma 2 yields for later increments: Instead of directly considering a counting constraint of the form “The player $EGO$ plays $a$ at most $k$ times out of $l$ of its own turns.”, we can start with “The player $EGO$ plays $\neg a$ at least lkl-k times out of lkl-k of its own turns.”, incrementally increase the constraint length to “The player $EGO$ plays $\neg a$ at least lkl-k times out of $l$ of its own turns.”, which translated back to the original counting constraint.
@@
Figure 4. Snippet of a situation graph for the game from Example 2. A state (“situation”) of the game graph is shown as a combination of a state from the original game graph (Figure 1) with a history that allows to check counting constraint fulfillment in the situation. For better readability, the history is written as table with the name of the counting constraints as header: $C_{1}=$ “The player $EGO$ plays $x$ at least $2$ times out of $4$ of its own turns.”, $C_{2}=$ “The player $EGO$ plays $y$ at least $1$ times out of $5$ of its own turns.”, $C_{3}=$ “The player $EGO$ plays $y$ at most $2$ times out of $3$ of its own turns.”, $C_{4}=$ “The player $ALTER$ plays $a\vee b$ at least $1$ time out of $2$ of its own turns.”.
@@
For each counting constraint $CC_{m}(p,a,k,l)$ with $m\in\{min,max\}$, $p\in\{EGO,ALTER\}$, $a$ being a logical formula over $(\Sigma_{p})$ as defined in Definition 2 , $\in\mathbb{N}, klk\leq l$ define a mapping
@@
[…]
@@
  • initial state being the situation $s_{init}=(s_{0},H_{init})$ with all entries in $H_{init}$ being nonenone,
@@
[…]
@@
[…]
@@
[…]
@@
Figure 5. Examples for situations being extensions of other situations. The notation is the same as in Example 3 and Figure 4.
@@
Figure 6. Overview of the incremental synthesis algorithm for a game with counting constraints $G=(A,\text{Win},CC)$ with increments over constraints of the form $CC_{min}(EGO,a,k,l)$
@@
  • Situations that violate any constraint of $ALTER$ are not added to the situation graph. For any situation that satisfies all constraints of player $A$ and is controlled by $ALTER$, ALTERALTER can choose an action as next turn, such that the reached situation also satisfies all of its counting constraints (as a consequence of the last property in Definition 2 and Theorem 1). Hence, all relevant behavior of $ALTER$ is still captured when omitting situations that violate $ALTER$-constraints.
@@
Figure 7. Illustration on how the situation graph is modified in the incremental synthesis over counting constraint length as depicted in Figure 6. Circles represent situations controlled by $EGO$, diamond-shaped situations are controlled by $ALTER$. Situations colored in gray represent extensions of situations in the winning region of a previous increment. Situations marked with doubled lines are considered to have a history that violates a constraint of $EGO$. Similarly, histories of situations with dashed lines are violating a constraint of $ALTER$.
@@
[…]
@@
[…]
@@
as coloring on $Sit_{\mathcal{C}}$. Situations (except for the new sink-locations) inherit their color from their state in $S$. Whenever a play enters the winning sink $W$, EGOEGO will win the play, since each situation in $W$ is visited infinitely often and is colored with 0, being the minimal coloring and even. Similarly, if a play enters the sink $L$, EGOEGO is going to loose, since $L$ will never be left, hence only finitely many visited situations are not in $L$ and all situations in $L$ are colored with the uneven color 1.
@@
  2. (2)
[…]